OpenStudy (anonymous):

[if \alpha and \beta are the zeroes of the given polynomialf(x)=5x ^{2}-7x+1,find the value of 1/\alpha and 1/\]

OpenStudy (anonymous):

parth help yaar plzzzzzzzz

Parth (parthkohli):

http://openstudy.com/study#/updates/4fed6d53e4b0bbec5cfcc8b0 Come here

OpenStudy (anonymous):

whts the full ques..?????

OpenStudy (anonymous):

so at the first post dude thats the question

OpenStudy (anonymous):

hey the other question is \[obtain all zeroes of 3x ^{4}+6x ^{3}-2x ^{2}-10x-5,if two of its zeroes are \sqrt{5}/\sqrt{3} and -\sqrt{5}/\sqrt{3}\]

OpenStudy (anonymous):

did u get

OpenStudy (anonymous):

oh ther is a problem in my tablet it is not displaying the special characters and equations

OpenStudy (anonymous):

k its fine allen it happens parth can u help me out plzzz

OpenStudy (anonymous):

did u find

OpenStudy (anonymous):

ur 1st ans is \[(7+\sqrt{29}) \div10\] and \[(7-\sqrt{29})\div10\]

OpenStudy (anonymous):

the values of \[1\div \alpha\] and \[1\div \beta \]